find the residue of f(z)=z^2+2z/(z+1)^2(z+4)at its poles


1
Expert's answer
2022-03-10T03:21:04-0500

The poles of "f(z)=\dfrac {z^2+2z}{(z+1)^2(z+4)}" are obtained from "(z+1)^2(z+4) = 0".


We get, "z = -1" is a pole of order 2 and "z = -4" is a simple pole.


"\begin{aligned}
Res\{f(z),-1\} &= \lim_{z=-1}\dfrac{d}{dz}\left\{(z+1)^2~f(z)\right\}\\
&=\lim_{z=-1}\dfrac{d}{dz}\left\{(z+1)^2 \dfrac{(z^2+2z)}{(z+1)^2(z+4)}\right\}\\
&=\lim_{z=-1}\dfrac{d}{dz}\left\{ \dfrac{z^2+2z}{z+4}\right\}\\
&=\lim_{z=-1} \left\{\dfrac{z^2+8z+8}{(z+4)^2}\right\}\\
&= \frac{1}{9}\\
Res\{f(z),-4\} &= \lim_{z=-4}\left\{(z+4)~f(z)\right\}\\
&=\lim_{z=-4}\left\{(z+4) \dfrac{(z^2+2z)}{(z+1)^2(z+4)}\right\}\\
&=\lim_{z=-4}\left\{ \dfrac{z^2+2z}{(z+1)^2}\right\}\\
&= 0
\end{aligned}"